### v2

# Maximal variation of linear systems

Abstract.
Let $X$ be a smooth projective variety over $\mathbb{C}$, and $L$ a line bundle on $X$. We say that the linear system $\lvert{L}\rvert$ has maximal variation if its elements have the maximum number $\dim\lvert{L}\rvert$ of moduli. We discuss four cases where this situation is expected: hypersurfaces, double coverings of $\mathbb{P}^{n}$, K3 surfaces, and abelian varieties.
Introduction
Let $X$ be a smooth projective variety over $\mathbb{C}$, and let $L$ be a line bundle on $X$. We assume that a general hypersurface $S$ in the linear system $\lvert{L}\rvert$ is smooth and connected. We say that $\lvert{L}\rvert$ has maximal variation if there are only finitely many $S^{\prime}$ in $\lvert{L}\rvert$ isomorphic to $S$. If there is a reasonable moduli space $\mathscr{M}$ for $S$ (this will be the case for most of our examples), this is equivalent to say that the natural rational map $\lvert{L}\rvert\dasharrow\mathscr{M}$ is generically finite, or equivalently that its image has dimension $\dim\lvert{L}\rvert$.
There is a simple differential criterion for this to happen, namely $H^{0}(T_{X|S})=0$ (§1). However this turns out to be surprisingly difficult to check. We will discuss four cases where maximal variation is expected: hypersurfaces (§2), double coverings of $\mathbb{P}^{n}$ (§3), K3 surfaces (§4), and abelian varieties (§5). We prove indeed that the variation is maximal in the latter case, while we have only partial results in the first three cases. Finally we discuss some more examples and counter-examples (§6 and 7).
1. A differential criterion
Proposition 1.
$1)$ If $H^{0}(S,T_{X|S})=0$, $\lvert{L}\rvert$ has maximal variation.
$2)$ If $H^{1}(X,\mathcal{O}_{X})=0$ and $H^{0}(S,T_{S})=0$, the converse holds.
Proof : Let $S$ be a general element of $\lvert{L}\rvert$, and let $k:\mathscr{S}\rightarrow B$ be a Kuranishi family for $S$. There is an open neighborhood $U$ of $[S]$ in $\lvert{L}\rvert$ and a map $\mu:U\rightarrow B$ such that the pull back of $k$ by $\mu$ is the family of hypersurfaces parameterized by $U$. $\lvert{L}\rvert$ has maximal variation if and only if $\mu$ is a local immersion at $[S]$, that is, the tangent map $T\mu$ is injective.
The exact sequence $0\rightarrow T_{S}\rightarrow T_{X|S}\rightarrow L_{|S}\rightarrow 0$ gives a cohomology exact sequence
$0\rightarrow H^{0}(T_{S})\rightarrow H^{0}(T_{X|S})\rightarrow H^{0}(L_{|S})\xrightarrow{\ \partial\ }H^{1}(T_{S})\,.$
Let $s$ be a section of $L$ with zero locus $S$. The exact sequence $0\rightarrow\mathcal{O}_{X}\xrightarrow{\ s\ }L\rightarrow L_{|S}\rightarrow 0$ gives a cohomology exact sequence
$0\rightarrow H^{0}(L)/(\mathbb{C}\!\cdot\!s)\xrightarrow{\ r\ }H^{0}(L_{|S})\rightarrow H^{1}(\mathcal{O}_{X})\,.$
The tangent space to $\lvert{L}\rvert$ at $S$ is $H^{0}(L)/(\mathbb{C}\!\cdot\!s)$, and the tangent space to $B$ at $\mu(S)$ is $H^{1}(T_{S})$; by [K-S, (12.4)], the tangent map to $\mu$ at $[S]$ is the composition
$T\mu:H^{0}(L)/(\mathbb{C}\!\cdot\!s)\xrightarrow{\ r\ }H^{0}(L_{|S})\xrightarrow{\ \partial\ }H^{1}(T_{S})\,.$
If $H^{0}(T_{X|S})=0$, $\partial$ is injective; since $r$ is injective, $T\mu$ is injective. Conversely if $T\mu$ is injective, under the hypotheses of 2) $r$ is an isomorphism, so $\partial$ is injective and $H^{0}(T_{X|S})=0$ since $H^{0}(T_{S})=0$.∎
2. Hypersurfaces
Let $X\subset\mathbb{P}^{n}$ be a smooth hypersurface of degree $d\geq 3$, defined by an equation $F=0$. We assume $n\geq 3$, and $d\geq 4$ if $n=3$. The Jacobian ring of $X$ is the graded artinian ring $R:=\mathbb{C}[x_{0},\ldots,x_{n}]/(F^{\prime}_{0},\ldots,F^{\prime}_{n})$, where we write $F^{\prime}_{i}=\dfrac{\partial F}{\partial x_{i}}\,\cdot$
Proposition 2.
The linear system $\lvert{\mathcal{O}_{X}(1)}\rvert$ has maximal variation if and only if the multiplication map $\times\ell:R_{d-1}\rightarrow R_{d}$ is injective for $\ell$ general in $R_{1}$.
The property that $\times\ell:R_{p-1}\rightarrow R_{p}$ is of maximal rank for $\ell$ general in $R_{1}$ and for all $p$ is called the weak Lefschetz property; it is conjectured that it holds for all rings of the form $\mathbb{C}[x_{0},\ldots,x_{n}]/(P_{0},\ldots,P_{n})$, with $P_{0},\ldots,P_{n}$ homogeneous with no common zeroes in $\mathbb{P}^{n}$ (see for instance [J-M]). This would imply that $\lvert{\mathcal{O}_{X}(1)}\rvert$ has always maximal variation. We have only a weaker result:
Corollary 1.
$\lvert{\mathcal{O}_{X}(1)}\rvert$ has maximal variation in the following cases:
$1)\ d\geq n+2$;
2)X2)\ X is a surface in $\mathbb{P}^{3}\ ($with $d\geq 4)$;
3)X3)\ X is a cubic threefold;
4)X4)\ X is general (among hypersurfaces of degree $d$ in $\mathbb{P}^{n}$).
Proof : 1) is proved in [B-M], improving by 1 the bound in [P-R-T]. This implies 2) for $d\geq 5$; for $d=4$ and $n=3$, the weak Lefschetz property is proved in [B-M-M-N].
3) is proved in [B2], as a consequence of [A-R].
4) The Fermat hypersurface $\sum X_{i}^{d}=0$ in $\mathbb{P}^{n}$ satisfies the weak Lefschetz property (in all degrees), because its Jacobian ring is isomorphic to the cohomology ring $H^{*}((\mathbb{P}^{d-2})^{n},\mathbb{C})$. Therefore the same holds for $X$ in a Zariski open subset of $\lvert{\mathcal{O}_{\mathbb{P}^{n}}(d)}\rvert$. ∎
Proof of the Proposition : We essentially repeat the proof of [B2], with one more detail for the case $n=3$.
Let $S$ be a smooth hyperplane section of $X$; we choose the coordinates so that it is defined by $x_{0}=0$. From the exact sequence
$0\rightarrow T_{X|S}\rightarrow T_{\mathbb{P}^{n}|S}\rightarrow\mathcal{O}_{S}(d)\rightarrow 0$
we see that $H^{0}(T_{X|S})$ is the kernel of the homomorphism $\varphi:H^{0}(T_{\mathbb{P}^{n}|S})\rightarrow H^{0}(\mathcal{O}_{S}(d))$.
From the commutative diagram of Euler exact sequences
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{0}(\mathcal{O}_{\mathbb{P}^{n}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\wr}$${H^{0}(\mathcal{O}_{\mathbb{P}^{n}}(1)^{n+1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{0}(T_{\mathbb{P}^{n}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{0}(\mathcal{O}_{S})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{0}(\mathcal{O}_{S}(1)^{n+1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{0}(T_{\mathbb{P}^{n}|S})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\mathcal{O}_{S})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{u}$${H^{1}(\mathcal{O}_{S}(1)^{n+1})}$
we get an exact sequence
$H^{0}(\mathbb{P}^{n},T_{\mathbb{P}^{n}})\xrightarrow{\ r\ }H^{0}(S,T_{\mathbb{P}^{n}|S})\rightarrow H^{1}(S,\mathcal{O}_{S})\xrightarrow{\ u\ }H^{1}(S,\mathcal{O}_{S}(1)^{n+1})\,.$
If $n\geq 4$ we have $H^{1}(S,\mathcal{O}_{S})=0$; if $n=3$ (so that $S$ is a curve), the homomorphism $u$ is the transpose of the map $H^{0}(S,\mathcal{O}_{S}(d-4))^{4}\xrightarrow{\ (x_{0},\ldots,x_{3})\ }H^{0}(S,\mathcal{O}_{S}(d-3))$, which is surjective since $d\geq 4$; thus $u$ is injective. In each case we conclude that the restriction map $r$ is surjective. It follows that $H^{0}(S,T_{\mathbb{P}^{n}|S})$ is generated by the elements $x_{i}\dfrac{\partial}{\partial x_{j}}$ for $i\geq 1$, $j\geq 0$, with the Euler relation $\sum\limits_{i\geq 1}x_{i}\dfrac{\partial}{\partial x_{i}}=0$.
Assume that $\times x_{0}:R_{d-1}\rightarrow R_{d}$ is injective. Let $V\in H^{0}(S,T_{\mathbb{P}^{n}|S})$; we can write $V=\sum\limits_{i}L_{i}\dfrac{\partial}{\partial x_{i}}$, where the $L_{i}$ are linear forms in $x_{1},\ldots,x_{n}$. The homomorphism $\varphi:H^{0}(T_{\mathbb{P}^{n}|S})\rightarrow H^{0}(\mathcal{O}_{S}(d))$ maps $V$ to $\sum L_{i}F^{\prime}_{i}$. If $\varphi(V)=0$ , there exists $G\in H^{0}(\mathbb{P}^{n},\mathcal{O}_{\mathbb{P}}(d-1))$ and $a\in\mathbb{C}$ such that $\sum L_{i}F^{\prime}_{i}=x_{0}G+aF$. Since $dF=\sum x_{i}F^{\prime}_{i}$, it follows that $x_{0}G$ is zero in $R_{d}$. By hypothesis this implies $G=0$ in $R_{d-1}$, that is, $G=\sum a_{i}F^{\prime}_{i}$ for some constants $a_{i}$. Therefore
$\sum_{i=0}^{n}(L_{i}-a_{i}x_{0}-bx_{i})F^{\prime}_{i}=0\qquad\mbox{with }b:=a/d\,.$
Since $(F^{\prime}_{0},\ldots,F^{\prime}_{n})$ is a regular sequence in $\mathbb{C}[x_{0},\ldots,x_{n}]$, this implies $L_{i}-a_{i}x_{0}-bx_{i}=0$ for each $i$. Since $x_{0}$ does not appear in $L_{i}$, we get :
$\bullet$ For $i\geq 1$, $a_{i}=0$ and $L_{i}=bx_{i}$;  $\bullet$ for $i=0$, $L_{0}=0$.
Hence $V=b\sum\limits_{i\geq 1}x_{i}\dfrac{\partial}{\partial x_{i}}=0$.
Conversely, suppose $H^{0}(S,T_{\mathbb{P}^{n}|S})=0$. Let $G\in H^{0}(\mathbb{P}^{n},\mathcal{O}_{\mathbb{P}}(d-1))$ such that $x_{0}G=0$ in $R_{d}$; there exist linear forms $L_{i}$ such that $x_{0}G=\sum L_{i}F^{\prime}_{i}$. Replacing $G$ by $G-\sum a_{i}F^{\prime}_{i}$, where $a_{i}$ is the coefficient of $x_{0}$ in $L_{i}$, we can assume that the $L_{i}$ are linear forms in $x_{1},\ldots,x_{n}$. Then the vector field $V=\sum L_{i}\dfrac{\partial}{\partial x_{i}}$ satisfies $\varphi(V)=0$, hence $V=0$. This means that there exists $c\in\mathbb{C}$ such that $L_{i}=cx_{i}$ for $i>0$ and $L_{0}=0$. Then $x_{0}G=c\sum\limits_{i>0}x_{i}F^{\prime}_{i}=c(dF-x_{0}F^{\prime}_{0})$, hence $cdF=x_{0}(G+cF^{\prime}_{0})$. Since $F$ is irreducible, this implies $c=0$, hence $G=0$. ∎
We now consider the linear system $\lvert{\mathcal{O}_{X}(e)}\rvert$ for $e\geq 2$.
Proposition 3.
$\lvert{\mathcal{O}_{X}(e)}\rvert$ has maximal variation in the following cases :
$1)\ \lvert{\mathcal{O}_{X}(1)}\rvert$ has maximal variation;
$2)\ e\geq d$.
(In particular, each of the conditions 1) to 3) of Corollary 1 implies that $\lvert{\mathcal{O}_{X}(e)}\rvert$ has maximal variation for all $e\geq 1$).
Proof : The proof of Proposition 2 works identically for the linear system $\lvert{\mathcal{O}_{X}(e)}\rvert$, replacing $x_{0}$ by a general form $h$ in $H^{0}(\mathcal{O}_{\mathbb{P}^{n}}(e))$. We find that $\lvert{\mathcal{O}_{X}(e)}\rvert$ has maximal variation if $\times h:R_{d-e}\rightarrow R_{d}$ is injective for $h$ general in $R_{e}$.
1) Let $\ell$ be a general element of $R_{1}$. If $\lvert{\mathcal{O}_{X}(1)}\rvert$ has maximal variation, $\times\ell:R_{d-1}\rightarrow R_{d}$ is injective (Proposition 2). Then $\times\ell:R_{p-1}\rightarrow R_{p}$ is injective for $p\leq d$ (if $\ell x=0$ for $x\in R_{p-1}$, we have $x\cdot R_{d-p}=0$, hence $x=0$). Therefore $\times\ell^{e}:R_{d-e}\rightarrow R_{d}$ is injective, and so is $\times h$ for $h$ general in $R_{e}$.
2) If $e>d$, we have $R_{d-e}=0$ so there is nothing to prove. If $e=d$, we just observe that $R_{d}\neq 0$, since the socle of $R$ is in degree $(n+1)(d-2)\geq d$.∎
3. Double coverings of $\mathbb{P}^{n}$
Let $\pi:X\rightarrow\mathbb{P}^{n}$ a double covering, branched along a smooth hypersurface $B\subset\mathbb{P}^{n}$ of (even) degree $d$. We assume $n\geq 2$, $d\geq 4$, and $d\geq 6$ if $n=2$. Let $F=0$ be an equation of $B$, and let $R=\mathbb{C}[x_{0},\ldots,x_{n}]/(F^{\prime}_{0},\ldots,F^{\prime}_{n})$ be the Jacobian ring of $F$.
Proposition 4.
The linear system $\lvert{\pi^{*}\mathcal{O}_{\mathbb{P}}(1)}\rvert$ has maximal variation if and only if the multiplication map $\times\ell:R_{d-1}\rightarrow R_{d}$ is injective for $\ell$ general in $R_{1}$.
Using again [B-M], we get:
Corollary 2.
If $d\geq n+2\ ($and $d\geq 6$ if $n=2)$, $\lvert{\pi^{*}\mathcal{O}_{\mathbb{P}}(1)}\rvert$ has maximal variation.
Proof of the Proposition : The proof is inspired by [D-H, Proposition 5.2], which treats the case $n=2$, $d=6$.
Put $\delta:=d/2$, and $\mathbb{P}=\mathbb{P}^{n}$. We have $\pi_{*}\Omega^{1}_{X}=\Omega^{1}_{\mathbb{P}}\oplus\Omega^{1}_{\mathbb{P}}\langle{B}\rangle(-\delta)$ [E-V, Lemma 3.16 (d)]. Using Grothendieck duality this gives
$\pi_{*}T_{X}=T_{\mathbb{P}}(-\delta)\oplus T_{\mathbb{P}}\langle{B}\rangle$
where $T_{\mathbb{P}}\langle{B}\rangle$ is the sheaf of vector fields on $\mathbb{P}$ tangent to $B$.
Let $H$ be a general hyperplane in $\mathbb{P}$, and $S:=\pi^{*}H$. We have
$\pi_{*}T_{X|S}=(\pi_{*}T_{X})_{|H}=T_{\mathbb{P}}(-\delta)_{|H}\oplus T_{\mathbb{P}}\langle{B}\rangle_{|H}\,$,
hence the vanishing of $H^{0}(T_{X|S})$ is equivalent to $H^{0}(T_{\mathbb{P}}(-\delta)_{|H})=H^{0}(T_{\mathbb{P}}\langle{B}\rangle_{|H})=0$.
1) Consider the Euler exact sequence restricted to $H$:
$0\rightarrow\mathcal{O}_{H}(-\delta)\rightarrow\mathcal{O}_{H}(-\delta+1)^{n+1}\rightarrow T_{\mathbb{P}}(-\delta)_{|H}\rightarrow 0\,.$
If $n\geq 3$, we have $H^{1}(\mathcal{O}_{H}(-\delta))=0$, hence $H^{0}(T_{\mathbb{P}}(-\delta)_{|H})=0$. If $n=2$ (so $H\cong\mathbb{P}^{1}$), the homomorphism $H^{1}(\mathcal{O}_{H}(-\delta))\rightarrow H^{1}(\mathcal{O}_{H}(-\delta+1)^{3})$ is the transpose of the map $H^{0}(\mathcal{O}_{H}(\delta-3)^{3})\xrightarrow{\ (x_{0},x_{1},x_{2})\ }H^{0}(\mathcal{O}_{H}(\delta-2))$, which is surjective since $\delta\geq 3$; hence again $H^{0}(T_{\mathbb{P}}(-\delta)_{|H})=0$.
2) We have an exact sequence
$0\rightarrow T_{\mathbb{P}}\langle{B}\rangle\rightarrow\mathcal{O}_{\mathbb{P}}(1)^{n+1}\xrightarrow{\ \varphi\ }\mathcal{O}_{\mathbb{P}}(d)\rightarrow 0\ ,\quad\mbox{where }\ \varphi(L_{0},\ldots,L_{n})=\sum L_{i}F^{\prime}_{i}\,.$
Therefore we have $H^{0}(T_{\mathbb{P}}\langle{B}\rangle_{|H})=0$ if and only if $\varphi_{|H}:H^{0}(\mathcal{O}_{H}(1)^{n+1})\rightarrow H^{0}(\mathcal{O}_{H}(d))$ is injective. Choose the coordinates so that $H$ is defined by $x_{0}=0$. Let $L_{0},\ldots,L_{n}$ be linear forms in $x_{1},\ldots,x_{n}$ such that $\varphi_{|H}(L_{0},\ldots,L_{n})=0$. This means that $\sum L_{i}F^{\prime}_{i}=x_{0}G$ for some $G\in H^{0}(\mathcal{O}_{\mathbb{P}}(d-1))$. If $\times x_{0}:R_{d-1}\rightarrow R_{d}$ is injective, we get $G=\sum a_{i}F^{\prime}_{i}$ for some scalars $a_{i}$, thus $\sum(L_{i}-a_{i}x_{0})F^{\prime}_{i}=0$. As before this implies $L_{i}=a_{i}x_{0}$, and since $x_{0}$ does not appear in $L_{i}$, $L_{i}=0$.
Conversely, suppose that $\varphi_{|H}$ is injective. Let $G\in R_{d-1}$ such that $x_{0}G=0$ in $R_{d}$: there exist linear forms $L_{0},\ldots,L_{n}$ such that $x_{0}G=\sum L_{i}F^{\prime}_{i}$. This implies that the restrictions $\bar{L}_{i}$ of $L_{i}$ to $H$ satisfy $\varphi_{|H}(\bar{L}_{0},\ldots,\bar{L}_{n})=0$, hence $\bar{L}_{i}=0$ for all $i$. Therefore $L_{i}=b_{i}x_{0}$ for some scalars $b_{i}$, so that $G=\sum b_{i}F^{\prime}_{i}$ is zero in $R_{d-1}$.∎
Remarks..- 1) Propositions 4 and 2 have the following curious consequence. The hypersurface $B$ embeds into $X$; under the hypotheses of Proposition 4, the linear system $\lvert{\pi^{*}\mathcal{O}_{\mathbb{P}}(1)}\rvert$ has maximal variation if and only if its restriction to $B$ has maximal variation. I don’t know a direct proof of this equivalence.
2) As in §2, the proof extends to the case of the linear system $\lvert{\pi^{*}\mathcal{O}_{\mathbb{P}}(e)}\rvert$ for $e>1$: we get that $\lvert{\pi^{*}\mathcal{O}_{\mathbb{P}}(e)}\rvert$ has maximal variation when $\lvert{\pi^{*}\mathcal{O}_{\mathbb{P}}(1)}\rvert$ does, or when $e\geq d$.
4. K3
For a polarized K3 surface $(X,L)$, it is conjectured that the linear system $\lvert{L}\rvert$ has maximal variation – in fact, it is shown in [D-H, Proposition 5.2] that this would follow from a conjecture of Matsushita on Lagrangian fibrations of hyperkähler manifolds. Here are some known cases:
Proposition 5.
Let $X$ be a K3 surface, and $L$ an ample line bundle on $X$, with $L^{2}=2g-2$. Then $\lvert{L}\rvert$ has maximal variation in the following cases:
$1)$ $g=2$, or $g=3$ and $(X,L)$ is not hyperelliptic;
$2)$ $g\geq 11$ and the Clifford index of the smooth curves in $\lvert{L}\rvert$ is $\geq 3$ ((this holds in particular when $\operatorname{Pic}(X)=\mathbb{Z}\cdot[L])$;
$3)$ $7\leq g\leq 10$ and $(X,L)$ is general;
$4)$ $L=M^{m}$ with $m\geq 2$, for an ample, globally generated line bundle $M$ such that the locus of non-reduced curves in $\lvert{H}\rvert$ has codimension $\geq 3$.
Proof : 1) The case $g=2$ is proved in [D-H, Proposition 5.5], see also Corollary 2. If $g=3$ and $(X,L)$ is not hyperelliptic, $X$ is a smooth quartic surface in $\mathbb{P}^{3}$ and $L=\mathcal{O}_{X}(1)$, so Corollary 1 applies.
2) is proved in [C-D-S, Proposition 2.4]; the last assertion follows from [L].
3) is proved in [C-D, Corollary 4.21], 4) in [D-H, Theorem 1.1].∎
5. Abelian varieties
Proposition 6.
Let $X$ be an abelian variety, and $L$ an ample line bundle on $X$. The linear system $\lvert{L}\rvert$ has maximal variation.
Proof : As in Proposition 1, we need to prove that the map $T\mu:H^{0}(L)/(\mathbb{C}\cdot s)\rightarrow H^{1}(T_{S})$ is injective. Let $s$ be a section of $L$ with zero locus $S$. The exact sequences
$0\rightarrow T_{S}\rightarrow T_{X|S}\rightarrow L_{|S}\rightarrow 0\qquad\mbox{and}\qquad 0\rightarrow\mathcal{O}_{X}\xrightarrow{\ s\ }L\rightarrow L_{|S}\rightarrow 0$
give rise to a diagram
${H^{0}(L)/(\mathbb{C}\!\cdot\!s)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r}$$\scriptstyle{T\mu}$${H^{0}(T_{X|S})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$$\scriptstyle{u}$${H^{0}(L_{|S})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\partial}$${H^{1}(T_{S})}$${H^{1}(\mathcal{O}_{X})}$
where the horizontal and vertical lines are exact.
[…]
Lemma 1.
$\tilde{u}:H^{0}(T_{X})\rightarrow H^{1}(\mathcal{O}_{X})$ is the cup-product with the Chern class $c_{1}(L)\in H^{1}(\Omega^{1}_{X})$.
Proof : We can find a finite covering $(U_{\alpha})_{\alpha\in I}$ of $X$ such that $L_{|U_{\alpha}}$ is free, generated by a section $e_{\alpha}$; on $U_{\alpha}\cap U_{\beta}$ we have $e_{\alpha}=g_{\alpha\beta}e_{\beta}$ for an invertible function $g_{\alpha\beta}$. We write $s_{|U_{\alpha}}=s_{\alpha}e_{\alpha}$ for a function $s_{\alpha}$ on $U_{\alpha}$; then $s_{\beta}=g_{\alpha\beta}s_{\alpha}$.
Let $V\in H^{0}(T_{X})$. Then $\varphi(V)$ is given on $U_{\alpha}\cap S$ by $(V\cdot s_{\alpha})e_{\alpha}$. This expression lifts to $U_{\alpha}$, and $u(V)=\partial\varphi(V)$ is the class of the Čech 1-cocycle $(\alpha,\beta)\mapsto c_{\alpha\beta}$ such that :
$c_{\alpha\beta}\,s=(V\cdot s_{\beta})e_{\beta}-(V\cdot s_{\alpha})e_{\alpha}=(V\cdot g_{\alpha\beta})g_{\alpha\beta}^{-1}\,s\ \mbox{ in }\,U_{\alpha}\cap U_{\beta}\,,\ \mbox{ hence }\ c_{\alpha\beta}=\langle{V,g_{\alpha\beta}^{-1}\,dg_{\alpha\beta}}\rangle\,.$
The class of the cocycle $(\alpha,\beta)\mapsto g_{\alpha\beta}^{-1}\,dg_{\alpha\beta}$ in $H^{1}(X,\Omega^{1}_{X})$ is the Chern class $c_{1}(L)$, hence the lemma.∎
Lemma 2.
$\tilde{u}$ is an isomorphism.
[…]
6. Some more examples
We collect here some other examples of linear systems with maximal variation. We always assume that the linear systems we consider contain a smooth connected curve.
$\bullet$ If $H^{0}(X,T_{X})=0$ and $L$ is ample, $\lvert{L^{m}}\rvert$ has maximal variation for $m\gg 0$. Indeed for any $S\in\lvert{L^{m}}\rvert$ there is an exact sequence $H^{0}(T_{X})\rightarrow H^{0}(T_{X|S})\rightarrow H^{1}(T_{X}\otimes L^{-m})$, and the right hand term vanishes for $m\gg 0$ since $L$ is ample.
$\bullet$ If $\Omega^{1}_{X}$ is ample, every linear system $\lvert{L}\rvert$ on $X$ has maximal variation. Indeed for any $S\in\lvert{L}\rvert$ the quotient $\Omega^{1}_{X|S}$ is ample, hence does not admit a nonzero map to $\mathcal{O}_{S}$.
For surfaces we have a more precise result:
Proposition 7.
Let $X$ be a surface such that $\Omega^{1}_{X}$ is generated by its global sections. Then any very ample linear system $\lvert{L}\rvert$ has maximal variation.
Proof : For a general 1-form $\omega\in H^{0}(\Omega^{1}_{X})$, the zero locus $Z(\omega)$ is finite, reduced and nonempty (since $c_{2}>0$). Let $p\in Z(\omega)$. Since $L$ is very ample, we can find a smooth curve $C\in\lvert{L}\rvert$ such that $C\cap Z(\omega)=\{p\}$. We have an exact sequence
$0\rightarrow K^{-1}_{X|C}(p)\rightarrow T_{X|C}\xrightarrow{\ \omega\ }\mathcal{O}_{C}(-p)\rightarrow 0\,;$
it suffices to prove $\deg(K^{-1}_{X|C}(p))<0$, that is, $K_{X}\cdot C>1$.
We now observe that $X$ is a minimal surface: if $R$ is a smooth curve in $X$, $\Omega^{1}_{R}$ is a quotient of $\Omega^{1}_{X}$, hence it must be globally generated, which implies $g(R)\geq 1$. Browsing through the classification of surfaces we get the following cases:
$\bullet\ X$ is an abelian surface. This case is settled by Proposition 6.
$\bullet\ X$ admits an elliptic fibration $p:X\rightarrow B$. Since $K_{X}$ is globally generated, it contains a line bundle $p^{*}M$ with $\deg M\geq 1$. On the other hand, a general curve $C\in\lvert{L}\rvert$ cannot be a fiber nor a section of $p$, so $K_{X}\cdot C\geq(p^{*}M)\cdot C>1$.
$\bullet\ X$ is of general type. Then by the index theorem we have $(K_{X}\cdot C)^{2}\geq K_{X}^{2}\cdot C^{2}>1$, hence $K_{X}\cdot C>1$ as required.∎
Proposition 8.
Let $X$ be a surface of general type such that $\operatorname{Pic}(X)=\mathbb{Z}\cdot[K_{X}]$ and that the Chern classes of $X$ satisfy $c_{1}^{2}>c_{2}$. The linear systems $\lvert{rK_{X}}\rvert$, for $r\geq 1$, have maximal variation.
Proof : The proof is borrowed from [B1]. Let $C$ be a smooth curve in $\lvert{rK_{X}}\rvert$. Suppose that $T_{X|C}$ has a nonzero section; this gives a surjective homomorphism $\Omega^{1}_{X}\rightarrow\mathcal{O}_{C}(-E)$, where $E$ is an effective divisor. The kernel $F$ of this homomorphism is a rank 2 vector bundle with
$c_{1}(F)=(r-1)c_{1}\ ,\quad c_{2}(F)=c_{2}-rc_{1}^{2}-\deg(E)\,$,
hence $c_{1}^{2}(F)-4c_{2}(F)\geq(r+1)^{2}c_{1}^{2}-4c_{2}\geq 4(c_{1}^{2}-c_{2})>0\,.$
By [Bo, §10] $F$ is unstable: there is an exact sequence
$0\rightarrow K_{X}^{a}\rightarrow F\rightarrow\mathscr{I}_{Z}K_{X}^{b}\rightarrow 0$
with $a\geq b$, and $Z$ a finite subscheme of $X$.
This gives $c_{1}(F)=-(a+b)c_{1}$, $c_{2}(F)=\deg(Z)+abc_{1}^{2}$, hence
$c_{1}^{2}(F)-4c_{2}(F)=(a-b)^{2}c_{1}^{2}-4\deg(Z)\,.$
Comparing with the previous expression gives
$(a-b)^{2}c_{1}^{2}\geq c_{1}(F)^{2}-4c_{2}(F)\geq(r+1)^{2}c_{1}^{2}-4c_{2}>(r^{2}+2r-3)c_{1}^{2}\geq(r-1)^{2}c_{1}^{2}$
Since $a\geq b$, we conclude ab>r1a-b>r-1, hence $a\geq 1$. This gives a nonzero homomorphism $K_{X}\rightarrow\Omega^{1}_{X}$, hence a nonzero section of $T_{X}$, which is impossible. ∎
Proposition 9.
Let $X=X_{1}\times X_{2}$, where $X_{1}$ and $X_{2}$ are smooth projective varieties of dimension $\geq 2$, and $H^{0}(X_{i},T_{X_{i}})=0$. Then any ample line bundle on $X$ has maximal variation.
Proof : Let $p_{i}$ be the projection of $X$ onto $X_{i}$. The restriction $\bar{p}_{i}:S\rightarrow X_{i}$ of $p_{i}$ to an ample divisor $S\subset X$ is surjective with connected fibers: indeed for $x\in X_{i}$, $\bar{p}_{i}^{-1}(x)=S\cap(\{x_{i}\}\times X_{1-i})$ is an ample divisor in $\{x_{i}\}\times X_{1-i}$, hence nonempty and connected. Therefore we have $(\bar{p}_{i})_{*}\mathcal{O}_{S}=\mathcal{O}_{X_{i}}$, and
$H^{0}((p_{i}^{*}T_{X_{i}})_{|S})=H^{0}(\bar{p}_{i}^{*}T_{X_{i}})=H^{0}(T_{X_{i}})=0\ ,\ \mbox{ hence }\ H^{0}(T_{X|S})=0\,.\qed$
7. Some counter-examples
$\bullet$ Suppose that a connected linear group $G$ acts non trivially on $X$. Then $G$ acts trivially on $\operatorname{Pic}(X)$, hence acts on each linear system $\lvert{L}\rvert$ on $X$. If this action is nontrivial, $\lvert{L}\rvert$ does not have maximal variation. This is the case for instance if $L$ is very ample, since one can find an element of $\lvert{L}\rvert$ passing through a point of $X$ and not containing its orbit.
$\bullet$ Let $C$ be a smooth projective curve, $E$ a rank 2 very ample vector bundle on $C$ (that is, $\mathcal{O}_{\mathbb{P}(E)}(1)$ is very ample), $X=\mathbb{P}(E)$, $L=\mathcal{O}_{\mathbb{P}(E)}(1)$. A general element of $\lvert{L}\rvert$ is isomorphic to $C$, so $\lvert{L}\rvert$ has no moduli.
$\bullet$ There are varieties of general type admitting a pencil with no moduli. Take for instance the Fermat hypersurface $F:\sum\limits_{i=0}^{n}x_{i}^{d}=0$. Consider the rational map $p:F\dasharrow\mathbb{P}^{1}$ given by $p(x_{0},\ldots,x_{n})=(x_{0},x_{1})$. Let $\hat{F}\rightarrow F$ be the blowing up of the codimension 2 subvariety given by $x_{0}=x_{1}=0$. Then $p$ extends to a morphism $\hat{p}:\hat{F}\rightarrow\mathbb{P}^{1}$ whose general fiber is isomorphic to the Fermat hypersurface of degree $d$ in $\mathbb{P}^{n-1}$. The corresponding linear system $\lvert{\hat{p}^{*}\mathcal{O}_{\mathbb{P}^{1}}(1)}\rvert$ has no moduli.
These examples lead to the following :
Question..- If $X$ is not uniruled, does every ample linear system $\lvert{L}\rvert$ have maximal variation?
It is probably overoptimistic to expect a positive answer, but it seems not obvious to find a counter-example.